OpenStudy (hannahc234):

What is the excluded value of the rational expression 2x + 6/ 4x - 8 ? A. -3 B. -2 C. 0 D. 2

OpenStudy (hannahc234):

I think it's d?

OpenStudy (xguardians):

-2, since it creates the denominator equal to zero.

OpenStudy (xguardians):

4x - 8 = ? 4(-2) - 8 = 0 You can't have zero in the denominator, it makes it irrational.

OpenStudy (stacey):

4(-2) - 8 = -8 - 8 = -16

OpenStudy (stacey):

The way that problem is written it appears that only 4x is in the denominator.

OpenStudy (stacey):

Is the problem supposed to look like this: \[2x+\frac{ 6 }{ 4x }-8\] or should there be parenthesis around 4x - 8 to make that binomial the denominator?

Directrix (directrix):

2x + 6 2(x + 3) --------- = -------- 4x - 8 4(x - 2) What value of x makes 4(x - 2) = 0 ? @HannahC234 That will be the excluded value.